# TTSA3R: Training-Free Temporal-Spatial Adaptive Persistent State for Streaming 3D Reconstruction

### v1

@@ Abstract
Streaming recurrent models enable efficient 3D reconstruction by maintaining persistent state representations. However, they suffer from catastrophic memory forgetting over long sequences due to balancing historical information with new observations. Recent methods alleviate this by deriving adaptive signals from attention perspective, but they operate on single dimensions without considering temporal and spatial consistency. To this end, we propose a training-free framework termed TTSA3R that leverages both temporal state evolution and spatial observation quality for adaptive state updates in 3D reconstruction. In particular, we devise a Temporal Adaptive Update Module that regulates update magnitude by analyzing temporal state evolution patterns. Then, a Spatial Contextual Update Module is introduced to localize spatial regions that require updates through observation-state alignment and scene dynamics. These complementary signals are finally fused to determine the state updating strategies. Extensive experiments demonstrate the effectiveness of TTSA3R in diverse 3D tasks. Moreover, our method exhibits only 15% error increase compared to over 200% degradation in baseline models on extended sequences, significantly improving long-term reconstruction stability. Our codes will be available soon.
@@ 1 Introduction
However, extended sequences reveal limited length generalization due to catastrophic forgetting [li2024incre, chen2025sab3r, chen2025ttt3r]. The uniform state update strategy [wang2025continuous] fails to prevent low-quality observations from overwriting historical information, which leads to accumulated errors and severe geometric distortions, as illustrated in Figure 1. In addition, spatial correspondence between states and observations is equally critical for determining update quality. To be specific, spatial interaction depends on both cross-attention alignment and feature consistency across frames. High cross-attention to regions with changing features indicates active geometric refinement, whereas cross-attention to regions with stable features suggests converged representations. Without jointly modeling these complementary signals, spatially irrelevant updates will corrupt stable geometry and necessary refinements may be missed. Recent training-free methods [chen2025ttt3r, shen2025mut3r] alleviate this issue through attention-based adaptive update mechanisms. However, these approaches only rely on single attention signals to guide state updates, which limits fine-grained state update control.
In this work, we present a training-free framework TTSA3R that mitigates the aforementioned issue through explicit temporal-spatial adaptive updates. We observe that state degradation stems from indiscriminate updates that fail to distinguish between stable geometry requiring preservation and outdated regions needing refinement. Therefore, effective updates should consider the analyses of both temporal and spatial signals. Specifically, temporal state evolution reveals the magnitude of change; spatial correspondence distinguishes whether high attention indicates active refinement with feature change or stable geometry with feature consistency. To this end, our approach introduces dedicated modules to analyze these complementary dimensions, thereby enabling fine-grained updates that preserve valuable information while integrating necessary observations.
@@ 3.1 Preliminaries: CUT3R Inference
Based on encoding results, a shared-weight ViT decoder [wein2022croco, wang2024dust3r] is used to process the current image tokens together with the previous global state and is the number of state tokens. The decoder outputs new state tokens and multi-layer features :
@@ 4.2 Camera Pose Estimation
We further evaluate 3D reconstruction on NRGBD [azinovic2022neural] dataset across different sequences. The quantitative results are shown in Figure 6. Compared to CUT3R [wang2025continuous] and TTT3R [chen2025ttt3r], our method shows better performance on all sequence lengths, which exhibits effective resistance to memory forgetting problem.
@@ 4.4 Ablation Studies and Analysis
This indicates that TAUM effectively mitigates error accumulation by adaptively controlling update magnitude based on temporal state evolution, which is particularly beneficial for maintaining trajectory consistency in pose estimation. Adding SCUM alone obtains 0.074 Abs Rel and 0.040 ATE, respectively. This shows that SCUM identifies spatial changing regions for updates based on observation-state alignment, preventing uniform updates across areas. When both modules are integrated into the framework, TTSA3R achieves the best performance with 0.064 Abs Rel and 0.026 ATE. This verifies that temporal-spatial adaptive updates effectively alleviate memory forgetting by selectively refining state representation.
Camera Trajectory Analysis. We visualize predicted camera trajectories on TUM-dynamics [sturm2012benchmark] dataset to analyse pose estimation. Figure 8 presents two representative sequences from indoor scenes with moving persons. It can be seen that the trajectories of CUT3R [wang2025continuous] reveal progressive drift accumulation during continuous motion and abrupt deviations during rapid viewpoint changes. Though recent approaches like TTT3R [chen2025ttt3r] reduce drift magnitude, noticeable deviations from the reference path remain. These observations reflect how existing recurrent methods lack effective adaptive mechanisms that respond to varying observations and temporal dynamics. Our predicted trajectories closely align with the reference path throughout the sequence.
@@ 5 Conclusions
In this work, we present a training-free 3D reconstruction framework TTSA3R that mitigates catastrophic memory forgetting of recurrent models for streaming reconstruction through an adaptive temporal-spatial mechanism. Specifically, the key contribution is a dual-module design that models temporal state dynamics and spatial correspondence, enabling complementary selective updates based on cross-frame evolution and observation quality. Consequently, this approach enables fine-grained control over memory persistence with real-time inference efficiency. Extensive experiments demonstrate that our method achieves competitive or superior performance compared to state-of-the-art pipelines across different benchmarks and tasks.

### v2

@@ Abstract
Streaming recurrent models enable efficient 3D reconstruction by maintaining persistent state representations. However, they suffer from catastrophic forgetting over long sequences due to balancing historical information with new observations. Recent methods alleviate this by deriving adaptive signals from attention perspective, but they operate on single dimensions without considering temporal and spatial consistency. To this end, we propose a training-free framework termed TTSA3R that leverages both temporal state evolution and spatial observation quality for adaptive state updates in 3D reconstruction. In particular, we devise a Temporal Adaptive Update Module that regulates update magnitude by analyzing temporal state evolution patterns. Then, a Spatial Contextual Update Module is introduced to localize spatial regions that require updates through observation-state alignment and scene dynamics. These complementary signals are finally fused to determine the state updating strategies. Extensive experiments demonstrate the effectiveness of TTSA3R in diverse 3D tasks. Moreover, our method exhibits only 1.33 error increase compared to over 4 degradation in the baseline model on extended sequences of 3D reconstruction, significantly improving long-term reconstruction stability. Our codes are available at https://github.com/anonus2357/ttsa3r.
@@ 1 Introduction
However, extended sequences reveal limited length generalization due to catastrophic forgetting [li2024incre, chen2025sab3r, chen2025ttt3r]. The uniform state update strategy [wang2025continuous] fails to prevent low-quality observations from overwriting historical information, which leads to accumulated errors and severe geometric distortions, as illustrated in Figure 1. In addition, spatial correspondence between states and observations is equally critical for determining update quality. To be specific, spatial interaction depends on both cross-attention alignment and feature consistency across frames. High cross-attention to regions with changing features indicates active geometric refinement, whereas cross-attention to regions with stable features suggests converged representations. Without jointly modeling these complementary signals, spatially irrelevant updates will corrupt stable geometry and necessary refinements may be missed. Recent methods [chen2025ttt3r, shen2025mut3r] alleviate this issue through attention-based adaptive update mechanisms. However, these approaches only rely on single attention signals to guide state updates, which limits fine-grained state update control.
In this work, we present a 3D reconstruction framework TTSA3R that mitigates the aforementioned issue through explicit temporal-spatial adaptive updates. We observe that state degradation stems from indiscriminate updates that fail to distinguish between stable geometry requiring preservation and outdated regions needing refinement. Therefore, effective updates should consider the analyses of both temporal and spatial signals. Specifically, temporal state evolution reveals the magnitude of change; spatial correspondence distinguishes whether high attention indicates active refinement with feature change or stable geometry with feature consistency. To this end, our method introduces dedicated modules to analyze these complementary dimensions, thus enabling fine-grained updates that preserve valuable information while integrating necessary observations.
@@ 3.1 Preliminaries: CUT3R Inference
Based on encoding results, a shared-weight transformer decoder [wein2022croco, wang2024dust3r] is used to process the current image tokens together with the previous global state and is the number of state tokens. The decoder outputs new state tokens and multi-layer features :
@@ 4.2 Camera Pose Estimation
We further evaluate 3D reconstruction on NRGBD [azinovic2022neural] dataset, as shown in Figure 6. Our method exhibits only 1.33x error increase compared to over 4x degradation in baseline CUT3R [wang2025continuous] as sequences extend from 50 to 250 frames. Compared to TTT3R [chen2025ttt3r], our method shows better performance on all sequence lengths, demonstrating more effective resistance to catastrophic forgetting.
@@ 4.4 Ablation Studies and Analysis
This indicates that TAUM effectively mitigates error accumulation by adaptively controlling update magnitude based on temporal state evolution, which is particularly beneficial for maintaining trajectory consistency in pose estimation. Adding SCUM alone obtains 0.074 Abs Rel and 0.040 ATE, respectively. This shows that SCUM identifies spatial changing regions for updates based on observation-state alignment, preventing uniform updates across areas. When both modules are integrated into the framework, TTSA3R achieves the best performance with 0.064 Abs Rel and 0.026 ATE. This verifies that temporal-spatial adaptive updates effectively alleviate catastrophic forgetting by selectively refining state representation.
Camera Trajectory Analysis. We visualize predicted camera trajectories on TUM-dynamics [sturm2012benchmark] dataset to analyze pose estimation. Figure 8 presents two representative sequences from indoor scenes with moving persons. It can be seen that the trajectories of CUT3R [wang2025continuous] reveal progressive drift accumulation during continuous motion and abrupt deviations during rapid viewpoint changes. Though recent approaches like TTT3R [chen2025ttt3r] reduce drift magnitude, noticeable deviations from the reference path remain. These observations reflect how existing recurrent methods lack effective adaptive mechanisms that respond to varying observations and temporal dynamics. Our predicted trajectories closely align with the reference path throughout the sequence.
@@ 5 Conclusions
In this work, we present a training-free 3D reconstruction framework TTSA3R that mitigates catastrophic forgetting of recurrent models for streaming reconstruction through an adaptive temporal-spatial mechanism. Specifically, the key contribution is a dual-module design that models temporal state dynamics and spatial correspondence, enabling complementary selective updates based on cross-frame evolution and observation quality. Consequently, this approach enables fine-grained control over memory persistence with real-time inference efficiency. Extensive experiments demonstrate that our method achieves competitive or superior performance compared to state-of-the-art pipelines across different benchmarks and tasks.